# Property prediction for ionic liquids without prior structural knowledge using limited experimental data: A data-driven neural recommender system leveraging transfer learning

### v1

@@ 2.2.2 Data sampling and COSMO simulations
After generating structural data for 2,268 cations and 311 anions, a stratified random sampling approach is applied. 50 anions are randomly selected for each cation, leading to a total of approximately 111,900 ionic liquids—about 15% of all 700,000+ possible cation-anion combinations considered in this study. We then run simulations to generate pre-training data for these ILs. COSMOtherm 2020 with TZVP parametrization is used to predict three properties: density (Preiss et al., 2009), viscosity (Eiden et al., 2011), and heat capacity (Preiss et al., 2009) at fixed temperature ($298\text{\,}\mathrm{K}$) and/or pressure ($1\text{\,}\mathrm{bar}$). Each property is computed using predefined models available in COSMO-RS, based on the parameter sets reported in the respective references.
@@ 2.2.3 Experimental datasets
For the fine-tuning step, experimental data for five IL properties are collected from published sources (Paduszynski, 2019a, b; Venkatraman et al., 2019; Low et al., 2020), comprising density (2,261 ILs), viscosity (1,995 ILs), surface tension (330 ILs), heat capacity (236 ILs), and melting point (2,178 ILs). After compiling all unique cations and anions across the datasets, data sparsity is evaluated. LABEL:fig:data_sparsity presents the total number of unique cations (2,268) and anions (311) included in the study, alongside the number of ionic liquids for which experimental data are available for each property. It is evident from the figure that the experimental datasets are highly sparse in comparison to the large number of possible cation-anion combinations. The total number of unique ILs across all properties is influenced by the overlap of cations and anions, with the highest coverage seen in density and melting point data, while surface tension and heat capacity exhibit significantly lower representation due to the limited number of unique ILs (330 and 236, respectively) compared to the total pool of 2,268 cations and 311 anions.
@@ 3.1 Model performance and transfer learning outcomes
The fine-tuned models are compared with the best-performing models previously reported in the literature. As shown in LABEL:tab:performance_comparison_1, the proposed framework achieves significant improvements for properties influenced by both molecular structure as well as temperature and pressure. For example, MAE improvements of 38% for density, 57% for viscosity, 75% for heat capacity, and 43% for surface tension are observed. These results confirm the benefit of the sequential learning approach used in this work. On the other hand, a decrease of 20% in model performance is noted for melting point, which depends only on molecular structure. This indicates that transfer learning may not offer as many benefits for structure-only properties. It is worth noting that the experimental dataset for viscosity, expressed as the natural logarithm of viscosity in mPa·s, ranges from 0 to 15. In contrast, the COSMO-RS simulated dataset at a fixed temperature of 298 K spans a wider range of 0 to 20, with some ionic liquids showing unrealistically high ln(viscosity) values of up to 80. These outliers, likely due to limitations in structure optimization or predictive models, indicate potential noise in the data, as they do not align well with other data during pre-training. Importantly, these noisy data points were intentionally retained to test our strategy’s ability to handle such noise, and it was found that they did not significantly affect the performance of the final fine-tuned model.
@@
The results demonstrate that the proposed transfer learning framework can successfully capture and transfer structural knowledge. This approach is particularly useful for predicting properties that depend on molecular structure as well as temperature and pressure. However, for properties dependent only on structure such as melting point, transferring such embeddings results in a decline in predictive performance compared to current benchmark. This may reflect the fact that the benchmark (Chen et al., 2023) utilizes more structural information compared to the proposed approach. The transfer learning framework effectively mitigates biases in experimental data, particularly for properties like surface tension and heat capacity, where experimental datasets include only a small subset of ionic liquids, ranging from 64 (surface tension) to 2,261 (density) ILs depending on the property. This approach enables property prediction across the entire combinatorial space of 705,348 ionic liquids formed from 2,268 cations and 311 anions, showcasing its ability to extrapolate with in a vast chemical spaces with limited experimental data. However, deviations in predictive accuracy are observed at the extreme ends of the heat capacity value range, occurring in both within-property and cross-property knowledge transfer during fine-tuning, as shown in LABEL:fig:density_transfer. This indicates that inter-property knowledge transfer is not the cause of the deviation, reinforcing the impact of sparse experimental data. Although the framework surpasses existing benchmarks for these properties, expanding experimental datasets for surface tension and heat capacity with additional ionic liquids is recommended to further enhance model performance and improve generalization across the chemical space.
The well-established interdependence of surface tension ($\sigma$), density ($\rho$), and viscosity ($\mu$) is supported by prior studies, which link these properties through molecular interactions and fluid characteristics (Sugden, 1924; Singh and Singh, 2011; Oroian, 2013). Sugden (1924) quantified the relationship between surface tension and density via the parachor $P$, a molecular descriptor reflecting chemical composition (see Eq. 1). Singh and Singh (2011) extended these correlations to ionic liquids, showing that denser fluids with stronger intermolecular forces exhibit higher viscosity and surface tension, often following empirical relations such as Eq. 2. Oroian (2013) further validated these interconnections in complex systems like honey, where temperature-dependent correlations indicate that viscosity and surface tension scale with density, as represented in Eq. 3. Our framework of transferring knowledge across properties essentially utilizes these correlations.
@@ 4 Conclusions
In this study, we integrate a neural recommender system in a transfer learning framework to predict multiple properties of ionic liquids using limited experimental data. The input features are learned sequentially: For molecular structures, we pre-train the NRS models using simulated COSMO data and then fine-tune them on limited experimental datasets to capture the effects of temperature and pressure. The results demonstrate that structural embeddings learned under fixed thermodynamic conditions can be effectively reused to predict properties influenced by both molecular structure as well as temperature and pressure. By efficiently generating COSMO data at fixed temperature and pressure to cover the entire cation-anion space, we ensure unbiased learning of structural features. These unbiased structural representations enable effective fine-tuning on biased experimental data across all temperatures and pressures.
Our approach achieves significant improvements in prediction accuracy for density, viscosity, heat capacity, and surface tension. These improvements are consistent across both within-property and cross-property transfer learning. For surface tension and melting point, for which no simulation is available, the framework successfully transfers knowledge from models pre-trained on other properties. However, no improvement is observed in the prediction of melting points, indicating that models with access to more structural information, such as Chen et al. (2023), achieve superior performance for properties which do not depend on temperature or pressure.
The proposed strategy enables property prediction across a vast chemical space of over 700,000 ILs, despite limited availability of experimental data. This highlights the framework’s scalability and strong extrapolation capabilities, offering a powerful tool for screening ILs in process design. We provide a ready-to-use Python function available on GitHub (link to be added) to directly input cation, anion, temperature, and/or pressure and predict the five properties included in this study.
Future extensions may include adapting the framework to predict additional thermophysical properties. For instance, interaction-based properties such as activity coefficients could be modeled using sequential knowledge transfer from activity coefficients at infinite dilution at fixed temperature to those at varying temperatures, and eventually to finite dilution data. This progression could enable prediction of complete vapor–liquid and liquid–liquid equilibra for solute–IL systems.
@@ Acknowledgments
The authors gratefully acknowledge the support of this research work from ERDF funds provided by the State of Saxony-Anhalt via the project “Further Development of the Center for Dynamic Systems (CDS) – Subproject MPI” (grant number: ZS/2023/12/182138). Sahil Sethi is also affiliated with the International Max Planck Research School for Systems and Process Engineering for a Sustainable Chemical Production (IMPRS SysProSus) at the Max Planck Institute for Dynamics of Complex Technical Systems Magdeburg. Special recognition is extended to Froze Jameel for his valuable assistance in configuring the COSMO calculations using TURBOMOLE. Additionally, the authors would like to thank Edgar Ivan Sanchez Medina for the constructive discussions concerning the methodology.
Data availability
The source code, trained models, and associated results from this study will be made available in a designated GitHub repository. The specific version utilized for this research will be specified upon publication.
Code availability
The models, along with training routines and experimental data, will be made publicly accessible on GitHub. The exact link will be provided upon publication.
@@
Sahil Sethi contributed to conceptualization, formal analysis, methodology, software development, and drafting the initial manuscript; Kai Sundmacher was responsible for funding acquisition, supervision, and manuscript review and editing; Caroline Ganzer offered conceptualization, methodology, supervision, and contributed to manuscript review and editing.
@@
Generative AI tools were used to assist in refining language and improving clarity in the manuscript. All scientific content, interpretations, and conclusions were developed by the authors.

### v2

@@ 2.2.2 Data sampling and COSMO simulations
After generating structural data for 2,268 cations and 311 anions, a stratified random sampling approach is applied. 50 anions are randomly selected for each cation, leading to a total of approximately 111,900 ionic liquids—about 15% of all 700,000+ possible cation-anion combinations considered in this study. We then run simulations to generate pre-training data for these ILs. COSMOtherm 2020 with TZVP parametrization is used to predict three properties: density (Preiss2009), viscosity (Eiden2011), and heat capacity (Preiss2009) at fixed temperature ($298\text{\,}\mathrm{K}$) and/or pressure ($1\text{\,}\mathrm{bar}$). Each property is computed using predefined models available in COSMO-RS, based on the parameter sets reported in the respective references.
@@ 2.2.3 Experimental datasets
For the fine-tuning step, experimental data for five IL properties are collected from published sources (Paduszynski2019a; Paduszynski2019b; Venkatraman2019; Low2020), concerning density (2,261 ILs), viscosity (1,995 ILs), surface tension (330 ILs), heat capacity (236 ILs), and melting point (2,178 ILs). After compiling all unique cations and anions across the datasets, data sparsity is evaluated. LABEL:fig:data_sparsity presents the total number of unique cations (2,268) and anions (311) included in the study, alongside the number of ionic liquids for which experimental data are available for each property. It is evident from the figure that the experimental datasets are highly sparse in comparison to the large number of possible cation-anion combinations. The total number of unique ILs across all properties is influenced by the overlap of cations and anions, with the highest coverage seen in density and melting point data, while surface tension and heat capacity exhibit significantly lower representation due to the limited number of unique ILs (330 and 236, respectively) compared to the total pool of 2,268 cations and 311 anions.
@@ 3.1 Model performance and transfer learning outcomes
The fine-tuned models are compared with the best-performing models previously reported in the literature. As shown in LABEL:tab:performance_comparison_1, the proposed framework achieves significant improvements for properties influenced by both molecular structure as well as temperature and pressure. For example, MAE improvements of 38% for density, 57% for viscosity, 75% for heat capacity, and 43% for surface tension are observed. These results confirm the benefit of the sequential learning approach proposed in this work. On the other hand, a decrease of 20% in model performance is observed for the melting point, which depends only on molecular structure. This indicates that transfer learning may not offer as many benefits for structure-only properties. It is worth noting that the experimental dataset for viscosity, expressed as the natural logarithm of viscosity in mPa·s, ranges from 0 to 15. In contrast, the COSMO-RS simulated dataset at a fixed temperature of 298 K spans a wider range of 0 to 0.02 Pa·s, with some ionic liquids showing unrealistically high ln(viscosity) values of up to 0.08 Pa·s. These outliers, likely due to limitations in structure optimization or predictive models, indicate potential noise in the data, as they do not align well with other data during pre-training. Importantly, these noisy data points were intentionally retained to test our strategy’s ability to handle such noise, and it was found that they did not significantly affect the performance of the final fine-tuned model.
@@
The results demonstrate that the proposed transfer learning framework can successfully capture and transfer structural knowledge. This approach is particularly useful for predicting properties that depend on molecular structure as well as temperature and pressure. However, for properties dependent only on structure such as melting point, transferring such embeddings results in a decline in predictive performance compared to current benchmark. This may reflect the fact that the benchmark (Chen2023) utilizes more structural information compared to the proposed approach. The transfer learning framework effectively mitigates biases in experimental data, particularly for properties like surface tension and heat capacity, where experimental datasets include only a small subset of ionic liquids, ranging from 64 (surface tension) to 2,261 (density) ILs depending on the property. This approach enables property prediction across the entire combinatorial space of 705,348 ionic liquids formed from 2,268 cations and 311 anions, showcasing its ability to extrapolate within a wide chemical space with limited experimental data. However, deviations in predictive accuracy are observed at the extreme ends of the heat capacity value range, occurring in both within-property and cross-property knowledge transfer during fine-tuning, as shown in LABEL:fig:density_transfer. This indicates that inter-property knowledge transfer is not the cause of the deviation, reinforcing the impact of sparse experimental data. Although the framework surpasses existing benchmarks for these properties, expanding experimental datasets for surface tension and heat capacity with additional ionic liquids is recommended to further enhance model performance and improve generalization across the chemical space.
The well-established interdependence of surface tension ($\sigma$), density ($\rho$), and dynamic viscosity ($\mu$) is supported by prior studies, which link these properties through molecular interactions and fluid characteristics (Sugden1924; Singh2011; Oroian2013). Sugden1924 quantified the relationship between surface tension and density via the parachor $P$, a molecular descriptor reflecting chemical composition (see Eq. 1). Singh2011 extended these correlations to ionic liquids, showing that denser fluids with stronger intermolecular forces exhibit higher viscosity and surface tension, often following empirical relations such as Eq. 2. Oroian2013 further validated these interconnections in complex substances like honey, where temperature-dependent correlations indicate that viscosity and surface tension scale with density, as represented in Eq. 3. Our framework of transferring knowledge across properties essentially utilizes these correlations.
@@ 4 Conclusions
In this study, we integrate a neural recommender system (NRS) in a transfer learning framework to predict multiple properties of ionic liquids using limited experimental data. The input features are learned sequentially: For molecular structures, we pre-train the NRS models using simulated COSMO data and then fine-tune them on limited experimental datasets to capture the effects of temperature and pressure. The results demonstrate that structural embeddings learned under fixed thermodynamic conditions can be effectively reused to predict properties influenced by both molecular structure as well as temperature and pressure. By efficiently generating COSMO data at fixed temperature and pressure to cover the entire cation-anion space, we ensure unbiased learning of structural features. These unbiased structural representations enable effective fine-tuning on biased experimental data across all temperatures and pressures.
Our approach achieves significant improvements in prediction accuracy for density, viscosity, heat capacity, and surface tension. These improvements are consistent across both within-property and cross-property transfer learning. For surface tension and melting point, for which no simulated data is available, the framework successfully transfers knowledge from models pre-trained on other properties. However, no improvement is observed in the prediction of melting points, indicating that models with access to more information on molecular structure, such as the one proposed by Chen2023, achieve superior performance for properties which do not depend on temperature or pressure.
The proposed strategy enables property prediction across a vast chemical space of over 700,000 ILs, despite limited availability of experimental data. This highlights the framework’s scalability and strong extrapolation capabilities, offering a powerful tool for screening ILs in process design. We provide a ready-to-use Python function available on GitHub (link to be added) to directly input the cation ID, anion ID, temperature, and/or pressure and predict the five properties included in this study.
Future extensions may include adapting the framework to predict additional thermophysical properties. For instance, interaction-based properties such as activity coefficients could be modeled using sequential knowledge transfer from activity coefficients at infinite dilution at fixed temperature to those at varying temperatures, and eventually to finite dilution data. This progression could enable prediction of complete vapor–liquid and liquid–liquid equilibria for solute–IL systems.
@@ Acknowledgments
The authors gratefully acknowledge the support of this research work by the European Regional Development Fund (ERDF), provided by the State of Saxony-Anhalt within the programme ”Research and Innovation”, via the project “Further Development of the Center for Dynamic Systems (CDS) – Subproject OVGU” (grant number: ZS/2023/12/182075). This work is also a part of the Research Cluster “SmartProSys: Intelligent Process Systems for the Sustainable Production of Chemicals” funded by the Ministry for Science, Energy, Climate Protection and the Environment of the State of Saxony-Anhalt in Germany. Sahil Sethi is also affiliated with the International Max Planck Research School for Systems and Process Engineering for a Sustainable Chemical Production (IMPRS SysProSus) at Otto von Guericke University Magdeburg. Special recognition is extended to Froze Jameel for his valuable assistance in configuring the COSMO calculations using TURBOMOLE. Additionally, the authors would like to thank their former colleague Edgar Ivan Sanchez Medina for the constructive discussions concerning the methodology.
Code and data availability
The source code, training routines, experimental data, trained models, and results from this study will be made publicly available in a designated GitHub repository. The link will be specified upon publication.
@@
Sahil Sethi contributed to conceptualization, formal analysis, methodology, software development, and drafting the initial manuscript; Kai Sundmacher was responsible for funding acquisition, supervision, and manuscript review and editing; Caroline Ganzer contributed to conceptualization, methodology, supervision, and manuscript review and editing.
@@
Generative AI tools were used to assist in refining language, editing codes, and improving clarity in the manuscript. All scientific content, interpretations, and conclusions were developed by the authors.